# Conformal Depression Prediction

### v1

@@ Abstract
While existing depression recognition methods based on deep learning show promise, their practical application is hindered by the lack of trustworthiness, as these deep models are often deployed as black box models, leaving us uncertain about the confidence of the model predictions. For high-risk clinical applications like depression recognition, uncertainty quantification is essential in decision-making. In this paper, we introduce conformal depression prediction (CDP), a depression recognition method with uncertainty quantification based on conformal prediction (CP), giving valid confidence intervals with theoretical coverage guarantees for the model predictions. CDP is a plug-and-play module that requires neither model retraining nor an assumption about the depression data distribution. As CDP provides only an average performance guarantee across all inputs rather than per-input performance guarantee, we propose CDP-ACC, an improved conformal prediction with approximate conditional coverage. CDP-ACC firstly estimates the prediction distribution through neighborhood relaxation, and then introduces a conformal score function by constructing nested sequences, so as to provide tighter prediction interval for each specific input. We empirically demonstrate the application of uncertainty quantification in depression recognition, and the effectiveness and superiority of CDP and CDP-ACC on the AVEC 2013 and AVEC 2014 datasets.
@@
depression recognition, uncertainty quantification, conformal prediction, approximate conditional coverage.
@@ I Introduction
Depression is a prevalent mental disorder presented as persistent feelings of sadness, debilitation, and loss of interest in activities [1, 2, 3]. Additionally, depression is often comorbid with many chronic medical diseases, with a substantially higher prevalence among patients affected by diseases such as cardiovascular, metabolic and neurological conditions compared to the general population [4]. It increases the risk of suicide and accounts for a substantial psychological burden [5, 6]. Therefore, AI researchers are working to develop automatic systems for diagnosing depression so that psychiatrists and psychologists can help patients in time. The existing primary diagnostic methods used involve mental health reports, such as the Beck Depression Inventory (BDI-II) [7] (target variable focused on in this paper), the Hamilton Depression Rating Scale (HRSD) [8], and the Patient Health Questionnaire (PHQ-8) [9]. The diagnostic process depends on interviews and demands considerable time and effort from both psychiatrists and patients. It heavily relies on the clinicians’ subjective experience, as well as the patients’ cognitive abilities and psychological states.
@@
Over the past decade, researchers have been exploring effective biomarkers and methods for recognizing depression. In the field of biomarkers, initial attention was focused on hand-crafted features such as Local Phase Quantization (LPQ) [10], Local Gabor Binary Patterns from Three Orthogonal Planes (LGBP-TOP) [11], and Local Binary Pattern from Three Orthogonal Planes (LBP-TOP) [12] extracted from facial videos. Other features explored include speech, facial action units (FAUs), facial landmarks, head poses, and gazes [13, 14]. The development of these hand-crafted features heavily relies on specific knowledge related to the depression recognition task. Acquiring such knowledge is time-consuming, subject to researchers’ subjective cognition and task specificity, lacking good generalization. Moreover, these features struggle to identify certain implicit and difficult-to-distinguish patterns of depression [15]. Fortunately, with the advent of deep learning [16], a new pathway has been opened for depression recognition tasks. Researchers can train end-to-end deep neural networks using depression-related data including facial video [17, 18, 19, 20], speech [21, 22, 13] and other sources. These deep neural networks can discover subtle, indistinguishable depression-related features and implicitly discriminate among them for improved prediction [23].
@@
Figure 1: The pipeline of uncertainty quantification for depression recognition. It can provide a confidence interval for depression prediction at any given confidence level, such as 90%.
While the depression recognition has yielded exciting results, leveraging the rapid development of deep learning and computer vision, its practicality and reliability have consistently raised concerns in clinical applications. Researchers are gradually realizing that deep learning models are black boxes [24, 25]. They lack interpretability, making it difficult to intuitively understand the recognized patterns. Additionally, they lack reliability, as the models may exhibit overconfidence in predicting failure, potentially leading to catastrophic risks. Similarly, depression recognition models also encounter these challenges.
@@
For depression recognition, the predictive performance should not be the only criterion determining the usefulness of a model. The lack of statistically rigorous uncertainty quantification is a key factor undermining the reliability of depression recognition [26]. Consider the application context of depression recognition. We are more willing to accept reliable predictions that meet strict statistical standards and provide insight into potential risks. Therefore, predictive performance should be considered in the context of principled uncertainty quantification. Beyond the depression prediction, we need additional outcomes to represent their uncertainty. It’s imperative to develop uncertainty quantification methods suitable for depression recognition. As shown in Figure 1, our aim is for the model to provide theoretically guaranteed and statistically rigorous confidence intervals in depression predictions. These intervals should cover the actual depression scores with any user-defined level of confidence.
@@
In this work, we propose conformal depression prediction (CDP) to quantify the uncertainty for depression recognition based on conformal prediction (CP [27, 28]). CDP can provide confidence intervals to satisfy marginal coverage for the actual depression scores. In addition, CDP as a plug-and-play module, can validly quantify uncertainty for most depression recognition models, without retraining the model. Furthermore, inspired by adaptive prediction sets (APS) in classification tasks, which can construct a prediction set for input adaptive to its prediction distribution, we propose conformal depression prediction with approximate conditional coverage (CDP-ACC) for providing adaptive confidence intervals for depression recognition to better uncertainty quantification.
@@
1. We develop a plug-and-play uncertainty quantification framework CDP for pre-trained depression recognition models. CDP can provide statistically rigorous confidence intervals that satisfy marginal coverage for most existing prediction models. To our knowledge, this is the first attempt to explore uncertainty quantification in depression recognition based on deep neural networks.
2. Furthermore, we propose CDP-ACC, an improved CDP method that satisfies approximate conditional coverage. Different from CDP, CDP-ACC estimates the prediction distribution through neighborhood relaxation, while a new conformal score is applied to ensure tight confidence intervals.
3. We conduct extensive experiments to demonstrate the application of uncertainty quantification in depression recognition and the effectiveness of our proposed method.
@@ II-A Depression Recognition
Over the past decade, the depression recognition models based on deep learning have attracted the interest of many researchers, giving rise to many meaningful researches. Researches focused on depression recognition using labeled depression scores from various modalities including speech, text, video, EEG signals, and the integration of these modalities in a multimodal way have taken a leading role.
@@
For single-modality depression recognition models, He et al. employed a Deep Convolutional Neural Network (DCNN) to extract depression-related features from speech data [21]. These learned features were subsequently integrated with additional hand-crafted features to predict depression scores. Zhao et al. introduced a hybrid network that integrates Self-Attention Networks (SAN) and DCNN to handle both low-level acoustic features and 3D Log-Mel spectrogram data [22]. Following this, feature fusion techniques along with support vector regression are employed for depression prediction. Melo et al. concentrated on efficiently extracting spatiotemporal features from facial videos associated with depression [17, 18]. They introduced a deep Multiscale Spatiotemporal Network (MSN) and a deep Maximization-Differentiation Network (MDN) to predict depression levels in facial videos. Sharma et al. proposed CNN-LSTM hybrid neural networks to extract depression features from EEG signals, which proved to be efficient for depression recognition [29]. Seal et al. developed the DeprNet, a DCNN designed for classification of EEG data from depressed and normal subjects [30].
@@
For multimodal depression recognition models, Niu et al. introduced a novel Spatio-Temporal Attention (STA) network and a Multimodal Attention Feature Fusion (MAFF) strategy to acquire a multimodal representation of depression cues from facial and speech spectrum [31]. Ahmed et al. leveraged features from text, speech, and facial landmarks for multimodal fusion [32]. They proposed an attention-based multimodal classifier with selective dropout and normalization. This approach simplifies the training process on a single neural network. Uddin et al. introduced a novel deep multi-modal framework that effectively integrates facial and verbal cues for automated depression recognition [33]. They extract features separately from video and speech and employ a Multi-modal Factorized Bilinear pooling (MFB) strategy to efficiently fuse these multi-modal features.
@@
These studies primarily concentrate on predicting the severity of depression and emphasize the accuracy of model predictions, while overlooking the analysis of uncertainty. Interestingly, works such as those of Melo et al. introduced minimizing expected error to learn distribution of depression levels [34], while Zhou et al. proposed DJ-LDML to learn depression representation using a label distribution and metric learning [20]. These studies implicitly explore uncertainty in depression recognition, yet without providing a clear and intuitive uncertainty quantification. To our best knowledge, this work is the first attempt to study the uncertainty quantification in depression recognition in modern contexts.
II-B Uncertainty in Affect Computing
@@
Uncertainty has been widely considered for various tasks in affective computing, such as emotion recognition, facial expression recognition, and apparent personality recognition [35, 36, 37, 38, 39, 40, 41]. In emotion recognition tasks, Harper et al. utilized a Bayesian framework to support emotional valence classification from the perspective of uncertainty [35]. Prabhu et al. proposed BNNs to model the label uncertainty based on subjectivity in emotion recognition [36]. Wu et al. employed deep evidence regression to jointly model aleatoric and epistemic uncertainties, aiming to enhance the performance of emotion attribute estimation [37]. Lo et al. utilized probabilistic uncertainty learning to extract information from low-resolution data and proposed the emotion wheel to learn label uncertainty, yielding robust facial expression recognition [38]. In facial expression recognition tasks, She et al. adopted a pairwise uncertainty estimation approach, assigning lower confidence scores to more ambiguous samples to address annotation ambiguities [39]. Le et al. proposed a method based on uncertainty-aware label distribution to adaptively construct the training sample distribution [40]. Additionally, Tellamekala et al. utilized a neural latent variable model to model aleatoric and epistemic uncertainty and integrated both to enhance the performance of apparent personality recognition [41].
@@
The uncertainty-inspired algorithms mentioned above are not suitable for depression recognition, as the available datasets with annotated BDI-II scores are limited. Re-stabilizing a well-performing, uncertainty-inspired depression recognition algorithm is not straightforward. Furthermore, while the above methods may achieve better performance in their respective tasks by introducing uncertainty-inspired heuristics, they do not establish a rigorously effective uncertainty quantification. In contrast to previous methods, our proposed CDP is a plug-and-play uncertainty quantification method for depression recognition, which does not require model retraining and does not assume the depression data distribution. Thanks to the conformal prediction [27, 42], our CDP can achieve valid confidence intervals with theoretical coverage guarantees for model predictions.
@@
However, current conformal prediction methods tend to produce unnecessarily conservative prediction intervals, resulting in the confidence interval to be unnecessarily wide [43]. In depression recognition, overly wide intervals are meaningless. For instance, the prediction for BDI-II (ranging from 0 to 64) has a 100% confidence interval of [0, 64], such a conclusion lacks practicality. Additionally, marginal coverage guarantees only valid average intervals, which overcovers simple subgroups and undercovers difficult ones [44]. Motivated by APS [45], we propose a conformal prediction method with approximate conditional coverage (CDP-ACC) that serves as a plug-and-play solution for the majority of pretrained regression-based depression predictor. Meanwhile, it can produce valid and adaptive confidence intervals based on the specific depression predictions.
@@ III-A Preliminaries
Formally, let us assume that we have the depression-associated data denoted by $X$ as input and the corresponding label denoted by $y$, which is the BDI-II score [7]. We randomly split the data into disjoint training set $\mathcal{D}_{train}$, calibration set $\mathcal{D}_{cal}=\{(X_{i},y_{i})\}_{i=1}^{n}$, and test set $\mathcal{D}_{test}$. We use $\mathcal{D}_{train}$ to train a DNN model $\hat{f}(\cdot)$ for depression recognition. Here, $\hat{y}\in\mathbb{R}$ denotes the prediction for $X$ by depression recognition model $\hat{f}(\cdot)$. We can formulate depression recognition as a vanilla regression task by minimizing the MSE loss (Eq. 1) as the learning objective.
$\begin{split}\mathcal{L}_{MSE}=\frac{1}{2N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}\end{split}$ (1)
@@
In depression recognition based on vanilla regression models, the depression prediction $\hat{y}$ is the expectation $\mathbb{E}[\mathcal{P}(y|X)]$ derived from estimating the conditional probability distribution $\mathcal{P}(y|X)$. However, We do not have access to the oracle of $\mathcal{P}(y|X)$. To quantify the uncertainty of depression prediction $\hat{y}$, we expect to construct an interval function $\mathcal{C}_{\alpha}(\cdot)$, which takes depression data and miscoverage rate $\alpha$ as input and outputs a confidence interval. The interval contains the depression target $y$ with any confidence level $1-\alpha$, as shown in Fig 1 ($\alpha=0.1$). Given the coverage of $1-\alpha$, $|\mathcal{C}(X_{n+1})|$ represents the width of the interval $X_{n+1}$, which indicates the uncertainty of $\hat{y}_{n+1}$. A wider interval $|\mathcal{C}(X_{n+1})|$ indicates greater uncertainty in the model’s depression prediction $\hat{y}_{n+1}$.
Figure 2: The pipeline of CDP. CDP requires a pre-trained depression recoginition model $\hat{f}$, a calibration set $\mathcal{D}_{cal}$ and a confidence level $1-\alpha$. As a post-processing method for uncertainty quantification, we utilize the depression predictions of the calibration set to calculate conformal scores. Subsequently, we employ the $(1-\alpha)$-th empirical quantile value $\hat{s}$ of conformal scores to form confidence intervals for $X_{test}$.
III-B Depression Recognition with UQ

### v3

@@ Abstract
While existing depression prediction methods based on deep learning show promise, their practical application is hindered by the lack of trustworthiness, as these deep models are often deployed as black box models, leaving us uncertain on the confidence of their predictions. For high-risk clinical applications like depression prediction, uncertainty quantification is essential in decision-making. In this paper, we introduce conformal depression prediction (CDP), a depression prediction method with uncertainty quantification based on conformal prediction (CP), giving valid confidence intervals with theoretical coverage guarantees for the model predictions. CDP is a plug-and-play module that requires neither model retraining nor an assumption about the depression data distribution. As CDP provides only an average coverage guarantee across all inputs rather than per-input performance guarantee, we further propose CDP-ACC, an improved conformal prediction with approximate conditional coverage. CDP-ACC firstly estimates the prediction distribution through neighborhood relaxation, and then introduces a conformal score function by constructing nested sequences, so as to provide a tighter prediction interval adaptive to specific input. We empirically demonstrate the application of CDP in uncertainty-aware facial depression prediction, as well as the effectiveness and superiority of CDP-ACC on the AVEC 2013 and AVEC 2014 datasets. Our code is publicly available at https://github.com/PushineLee/CDP.
@@
facial depression prediction, uncertainty quantification, conformal prediction, approximate conditional coverage.
@@ I Introduction
Depression is a prevalent mental disorder presented as persistent feelings of sadness, debilitation, and loss of interest in activities [1, 2, 3]. It increases the risk of suicide and accounts for a substantial psychological burden [4, 5]. The existing primary diagnostic methods used involve mental health reports, such as the Beck Depression Inventory (BDI-II) [6] (target variable focused on in this paper), the Hamilton Depression Rating Scale (HRSD) [7], and the Patient Health Questionnaire (PHQ-8) [8]. The diagnostic process depends on interviews and demands considerable time and effort from both psychiatrists and patients. Also, it heavily relies on the clinicians’ subjective experience, as well as the patients’ cognitive abilities and psychological states.
@@
Over the past decade, continuous efforts have been made to explore effective biomarkers and methods for automated depression prediction. For depression biomarkers, initial attention was focused on hand-crafted features such as Local Binary Patterns (LBP) [9], Local Gabor Binary Patterns from Three Orthogonal Planes (LGBP-TOP) [10], and Local Binary Pattern from Three Orthogonal Planes (LBP-TOP) [11, 12] extracted from facial videos. Handcrafted features extracted from other behavioral signals have also been employed in this task, including speech, facial action units, facial landmarks, head poses, and gazes [13, 14]. The development of these hand-crafted features heavily relies on specific knowledge related to the depression prediction tasks. Acquiring such knowledge is time-consuming, subject to researchers’ subjective cognition and task specificity, and lacks good generalization. Moreover, these features struggle to identify certain implicit and difficult-to-distinguish patterns of depression [15]. Fortunately, with the advent of deep learning [16], a new pathway has been opened for depression prediction tasks. In this pipeline end-to-end deep neural networks can be trained using depression data including facial video [17, 18, 19, 20], speech [21, 22, 13] and other data sources. These deep models are expected to discover subtle, indistinguishable depression-related features and implicitly discriminate among them for improved prediction.
@@
Figure 1: Uncertainty quantification for facial depression prediction can provide an additional confidence interval to score prediction at any given confidence level, such as 90% [^1] To protect personal privacy, we show facial landmarks instead of the original facial images..
@@
While the rapid development of deep learning and computer vision has led to exciting progress in depression prediction, the reliability and trustworthiness of deep learning models-considered black boxes [23, 24, 25]-have not been well addressed. These models may exhibit overconfidence in predicting failures, potentially leading to catastrophic risks. For automated depression prediction, the accuracy performance should not be the only criterion determining the usefulness of a predictive model. The lack of statistically rigorous uncertainty quantification is a key factor undermining the reliability of depression prediction. Consider the application context of depression prediction. We are more willing to accept reliable predictions that provide insight into potential risks. Beyond the depression severity prediction, we need additional outcomes to represent their uncertainty. As shown in Fig. 1, our aim is to provide statistically rigorous confidence intervals in depression prediction. These intervals should cover the actual depression scores with any user-defined level of confidence.
@@
In this work, we propose conformal depression prediction (CDP) to quantify the uncertainty of depression severity estimation based on conformal prediction (CP [26, 27]). CDP can provide confidence intervals to satisfy marginal coverage for the actual depression scores. We show that, CDP as a plug-and-play module, can validly quantify uncertainty for most depression prediction models, without retraining the model. Furthermore, inspired by the adaptive prediction sets (APS) [28] in uncertainty-aware classification tasks, which construct a prediction set per-input adaptive to its prediction distribution, we propose conformal depression prediction with approximate conditional coverage (CDP-ACC) to overcome the shortcomings of conformal prediction with marginal coverage, providing adaptive confidence intervals for depression prediction.
@@
1. We propose an uncertainty quantification framework CDP for facial depression prediction. CDP provides statistically rigorous confidence intervals that satisfy marginal coverage for most existing facial depression prediction models.
2. We develop CDP-ACC, an improved CDP algorithm that satisfies approximate conditional coverage, yielding tighter confidence intervals while maintaining good coverage.
3. We conduct extensive experiments to demonstrate the effectiveness of our proposed method in uncertainty quantification of facial depression prediction.
@@ II-A Depression Prediction
Over the past decade, depression prediction based on deep learning has attracted significant interest and attention in the field, leading to many meaningful research advancements. Studies using labeled depression data from various modalities, including speech, text, video, and EEG signals, as well as the integration of these modalities, have taken a leading role.
@@
For single-modality depression prediction, He et al. employed a Deep Convolutional Neural Network (DCNN) to extract depression-related features from speech data [21]. These learned features were subsequently integrated with additional hand-crafted features to predict depression scores. Zhao et al. introduced a hybrid network that integrates Self-Attention Networks (SAN) and DCNN to handle both low-level acoustic features and 3D Log-Mel spectrogram data [22]. Following this, feature fusion techniques along with support vector regression are employed for depression prediction. Melo et al. concentrated on efficiently extracting spatiotemporal features from facial videos associated with depression [17, 18]. They introduced a deep Multiscale Spatiotemporal Network (MSN) and a deep Maximization-Differentiation Network (MDN) to predict depression levels. Sharma et al. proposed CNN-LSTM hybrid networks to extract depression features from EEG signals [29]. Seal et al. developed the DeprNet, a DCNN designed for classification of EEG data from depressed and normal subjects [30].
@@
For multimodal depression prediction, Niu et al. introduced a novel Spatio-Temporal Attention (STA) network and a Multimodal Attention Feature Fusion (MAFF) strategy to acquire a multimodal representation of depression cues from facial and speech spectrum [31]. Uddin et al. introduced a novel deep multi-modal framework that effectively integrates facial and verbal cues for automated depression prediction [32]. They extracted features separately from video and speech and employed a Multi-modal Factorized Bilinear pooling (MFB) strategy to efficiently fuse these multi-modal features. Niu et al. proposed a novel multi-layer perceptron (MLP) architecture named DepressionMLP, designed to automatically predict depression levels through facial keypoints and action units (FKRS and AURS) [33].
@@
These studies, either using single-modality data or multimodal data, primarily focus on the prediction accuracy, while overlooking the analysis of predictive uncertainty. Interestingly, works such as those of Melo et al. introduced minimizing expected error to learn distribution of depression levels [34], while Zhou et al. proposed DJ-LDML to learn depression representation using a label distribution and metric learning [20]. These studies implicitly explore uncertainty in depression prediction, yet without providing a clear and intuitive uncertainty quantification. The method closely related to ours is the one proposed by Nouretdinov et al. [35], which leveraged conformal prediction for uncertainty quantification of MRI-based depression classification. Different from the work [35] that dedicated to the classification task, we focus on regression task in the context of deep learning, and we further develop an approximate conditional coverage (ACC) algorithm to overcome the shortcomings of conformal prediction with marginal coverage.
II-B Uncertainty Estimation in Affective Computing
@@
Uncertainty has been widely considered for various tasks in affective computing, such as emotion recognition, facial expression recognition, and apparent personality recognition [36, 37, 38, 39, 40, 41, 42]. In emotion recognition, Harper et al. utilized a Bayesian framework to support emotional valence classification from the perspective of uncertainty [36]. Prabhu et al. proposed BNNs to model the label uncertainty based on subjectivity in emotion recognition [37]. Wu et al. employed deep evidence regression to jointly model aleatoric and epistemic uncertainties, aiming to enhance the performance of emotion attribute estimation [38]. Tellamekala et al. introduced calibrated and ordinal latent distribution to capture modality-wise uncertainty for multimodal fusion in emotion recognition [43]. Lo et al. utilized probabilistic uncertainty learning to extract information from low-resolution data and proposed the emotion wheel to learn label uncertainty, yielding robust facial expression recognition [39]. In facial expression recognition, She et al. adopted a pairwise uncertainty estimation approach, assigning lower confidence scores to more ambiguous samples to address annotation ambiguities [40]. Le et al. proposed a method based on uncertainty-aware label distribution to adaptively construct the training sample distribution [41]. Additionally, Tellamekala et al. utilized a neural latent variable model to model aleatoric and epistemic uncertainty and integrated both to enhance the performance of apparent personality recognition [42].
@@
The uncertainty-inspired algorithms mentioned above may not be suitable for depression prediction, as the available datasets with annotated depression scores are limited. Re-stabilizing a well-performing, uncertainty-inspired depression prediction algorithm is not straightforward. Furthermore, while the above methods may achieve better accuracy performance in their respective tasks by introducing uncertainty-inspired heuristics, they do not establish a rigorously valid uncertainty quantification. In contrast to previous methods, our proposed CDP is a plug-and-play uncertainty quantification method for depression prediction, which does not require model retraining and does not assume the depression data distribution. Thanks to the conformal prediction [26, 44], our CDP can achieve valid confidence intervals with theoretical coverage guarantees for depression predictions.
Figure 2: The pipeline of CDP. CDP takes as input a pre-trained depression prediction model $\hat{f}$, a calibration set $\mathcal{D}_{cal}$ and a given confidence level $1-\alpha$. CDP sets the prediction biases (on the calibration set) as the conformal scores and employs the $(1-\alpha)$-th empirical quantile of the conformal scores to form confidence interval for the prediction $\hat{f}(X_{test})$.
@@
On the other hand, current conformal prediction methods tend to produce unnecessarily conservative intervals, resulting in the confidence interval to be unnecessarily wide [45]. In depression prediction, overly wide intervals are meaningless. For instance, the prediction with BDI-II (ranging from 0 to 63) has a 100% confidence interval of [0, 63]. Clearly, such a predictive uncertainty lacks practicality. Additionally, marginal coverage guarantees only valid average intervals, which overcovers simple subgroups and undercovers difficult ones [46]. Differently, our proposed conformal prediction method with approximate conditional coverage (CDP-ACC) can produce valid and tighter confidence interval adaptive to the specific depression prediction.
@@ III-A Preliminaries
Formally, let us assume that we have the facial depression-associated video data denoted by $X$ as input and the corresponding label denoted by $y$ (the BDI-II score [6], for example). We randomly split the data into disjoint training set $\mathcal{D}_{train}$, calibration set $\mathcal{D}_{cal}=\{(X_{i},y_{i})\}_{i=1}^{N}$, and test set $\mathcal{D}_{test}$. We use $\mathcal{D}_{train}$ to train a DNN model $\hat{f}(\cdot)$ for depression prediction. Here, $\hat{y}\in\mathbb{R}$ denotes the point prediction for $X$ by $\hat{f}(\cdot)$. In the deep learning context, we can formulate the depression prediction as a deep regression task by minimizing the mean squared error (MSE) loss as the learning objective.
$\begin{split}\mathcal{L}_{MSE}=\frac{1}{2|\mathcal{D}_{train}|}\sum_{i=1}^{|\mathcal{D}_{train}|}(\hat{y}_{i}-y_{i})^{2}\end{split}$ (1)
where $|\mathcal{D}_{train}|$ denotes the number of data in $\mathcal{D}_{train}$.
@@
In depression prediction based on such regression models, the point prediction $\hat{y}$ is the expectation $\mathbb{E}[y|X]$ derived from estimating the conditional probability distribution $P(y|X)$. However, we do not have access to the oracle of $P(y|X)$. To quantify the uncertainty of depression prediction $\hat{y}$, we expect to construct an interval function $\mathcal{C}_{\alpha}(X)$, which takes depression data $X$ and miscoverage rate $\alpha$ as input and outputs a confidence interval. The interval contains the depression target $y$ with any confidence level $1-\alpha$, as shown in Fig. 1 ($\alpha=0.1$). Clearly, wider interval $|\mathcal{C}_{\alpha}(X)|$ indicates greater uncertainty for the model’s prediction $\hat{y}$.
III-B Depression Prediction with UQ